# Rethinking Gating Mechanism in Sparse MoE: Handling Arbitrary Modality Inputs with Confidence-Guided Gate

### v1

@@ 3.2 Missing Modality Modeling: Two-Stage Imputation
Inspired by the modality gap phenomenon observed in contrastive learning [19], we hypothesize that a similar gap exists in our multimodal setting—and that it may even contribute positively to model performance, as illustrated in Appendix Figure 4. The presence of this modality gap suggests that imputing a missing modality using only shared information from the remaining modalities is inherently challenging, due to the distinct representational spaces each modality occupies. Ideally, the imputed modality should preserve the characteristics of its original distribution without being significantly biased by other modalities. To address this, we propose a Two-stage Imputation strategy as following:
Pre-Imputation: When a modality $M_{i}\in\mathbb{R}^{d}$ is missing, we generate an pre-imputed modality embedding $\bar{M}_{i}$ by sampling $K$ instances from the modality pool $M_{i,n}=[M_{i,o},M_{i,1},\cdots,M_{i,n}]$, where $M_{i,n}$ represent the $i$-th modality from $n$-th instance modality pool. The modality pool contains all existing modality data for given number of modalities in training set. To remove the instance-specific feature and preserve the modality-common feature, we take an average over the $K$ instances as pre-imputed feature. The randomized selection of instances from the training set further aids in eliminating instance-specific noise during backpropagation. Thus, we send the existing and pre-imputed modality into backbone model to extract interaction between modality. Naturally, increasing $K$ yields more representative modality features but reduces stochasticity, potentially causing all imputed modalities to converge to the same representation across different missing-modality samples and the interaction between other modalities may be more deterministic and uninformative. We consider this $K$ as one of hyperparameter to finetune.
$\bar{M}_{i}=\frac{1}{K}\sum^{K}_{n=1}\mathbf{1}M_{i,n}$ (6)
@@
Where $\mathbf{1}\in\{0,1\}^{n}$ is the selection vector for $M_{i}$ and $\mathbf{1}_{i}=1$ represents the $i$-th instance from modality pool is selected.
@@
Post-Imputation: To add up instance-specific information from other modality, we propose a post-modality imputation to further capture the interactions after the MoE forward. Unlike conventional approaches that solely rely on the fused representation from one network, our method leverages information from multiple experts networks by selectively incorporating tokens beyond the Top-K experts. This allows the model to absorb nuanced, instance-specific features and interaction between modalities from different views of expert. To implement this, we employ a sparse attention mechanism, under the assumption that only a subset of modality-specific information is meaningfully shared across existing modalities. By enforcing sparsity, the model is encouraged to attend only to relevant parts of the available modalities, thereby reducing the risk of injecting modality-irrelevant or noisy information into the imputed representation, especially in heterogeneous or weakly correlated multimodal settings. Specifically, given a missing modality $M_{i}\in\mathbb{R}^{s\times d}$ and existing modality $M_{j}\in\mathbb{R}^{s\times d}$, where $s$ is the sequence length and $d$ is the hidden dimension. Each token of existing modality $t_{s}\in\mathbb{R}^{d}$ is sent to Top-K different experts to generate $K$ specialized tokens $t_{s,e}\in\mathbb{R}^{d}$, where $s$ represents the sequential position and $e$ represents expert index. We concatenate the specialized tokens $t_{s,e}$ in sequence dimension and apply cross attention operation over those specialized tokens and select only Top-K correlated specialized tokens for $\bar{M}_{i}$ to generate final imputed modality $M_{i}^{*}$.
$M_{i}^{*}=\bar{M}_{i}+SparseCrossAttention(\bar{M}_{i},M_{j},M_{j})$ (7)
We set the number of tokens $K$ selected from $t_{s,e}$ as $K=\frac{s(|M|-1)}{B}$, where $|M|$ represents the number of modality in dataset and $B$ is a sparsity hyperparameter, we set $B$ as 4 by default.
@@ 4 Experiment
4.1 Dataset
MIMIC-III [16] is a public dataset contains irregular time series and clinical notes modality. We follow the preprocessing scripts from MMIMI-III benchmark [13] and [38] to form multimodal patient instance. However, [38] filter the instances with missing modality and we modified the scripts to preserve with instances with missing modality. MIMIC-IV [17] is a public dataset contains irregular time series, clinical notes, ECG signal and chest image. We follow the preprocessing scripts from MedFuse [14] to preprocess irregular time series and further preprocess clinical notes, ECG signal and chest image by ourself. We follow the dataset split in MedFuse [14] and form multimodal patient instance. CMU-MOSI [36] is a multimodal sentiment analysis dataset comprising audio, video, and text modalities. CMU-MOSEI [2] is an extension of the CMU-MOSI dataset, utilizing the same modalities—audio, text, and video from YouTube recordings. We follow preprocessing script of CMU-MOSI and CMU-MOSEI from MMML[22].
4.2 Baselines
SMIL[20] impute missing modality with Bayesian meta-learning. ShaSpec[29] leverages multiple loss function to learn the modality-share and modality-specific information and impute the missing modality with modality-shared information. TF[37] extract the modality interaction by multimodal sub-networks and tensor fusion layer. mmFormer[28] is a typical multimodal transformer with multihead attention mechanism. LlMoE[22] achieves the multimodal learning with contrastive learning. FuseMoE[12] respect the missing modality and jointly learn single and missing modality combination data input and FuseMoE implement Laplacian gating mechanism for modality fusoin and expert selection. In experiment, we denote FuseMoE-S as softmax gating and FuseMoE-L as Laplacian gating. FlexMoE[35] assigns the combination of modality to different experts and impute the missing modality by modality bank. For our approach, we use ConfSMoE-T denotes Token-Level confidence fusion and ConfSMoE-E denotes Expert-level confidence fusion.
@@ 4.3 Experiment Setting
To ensure a fair comparison with baseline methods, we adopt the best-performing hyperparameter settings reported in their respective papers and official implementations. Additionally, we standardize the hidden dimension to 128 across all models. To eliminate confounding effects introduced by varying encoder architectures, we use the same modality-specific encoders for all models, as differences in encoder design can significantly influence final performance. For the MIMIC-III and MIMIC-IV datasets, we employ a pretrained ClinicalBERT[1] for textual data, a one-layer patch embedding module for both irregular time series and ECG signals, and a pretrained DenseNet[7] from TorchXRayVision for chest X-ray (CXR) images. Experiment Setting I: Natural Missing Modality. This setting reflects real-world missingness patterns observed in clinical datasets. Experiment Setting II: Random Modality Dropout. A fixed percentage of modalities are randomly dropped for each instance, while ensuring that each instance retains at least one modality in both train and test set. Experiment Setting III: Asymmetric Modality Dropout. Half of the modalities are randomly dropped during training, while testing is performed with only one or two modalities present per instance, simulating domain shifts or deployment constraints. All experiments are 3-Fold cross-validate in different seeds for fair comparison.
@@ 4.4 Primary Results
Table 2: Experiment setting II: Main results on CMU-MOSI
Missing Rate Metric ShaSpec TF mmFormer LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfMoE-T ConfMoE-E
50% F1 36.39 $\pm$ 0.97 38.46 $\pm$ 1.09 41.14 $\pm$ 0.31 41.54 $\pm$ 0.77 38.55 $\pm$ 0.34 40.90 $\pm$ 0.49 41.95 $\pm$ 0.48 43.89 $\pm$ 0.51 42.48 $\pm$ 0.12
AUC 55.60 $\pm$ 0.12 60.49 $\pm$ 0.80 61.87 $\pm$ 0.67 67.43 $\pm$ 0.34 62.28 $\pm$ 0.28 65.73 $\pm$ 0.30 66.47 $\pm$ 0.49 71.51 $\pm$ 0.53 69.04 $\pm$ 0.32
Experiment Setting II: In Table 2 and Appendix E Table 10, we implement Random Modality Dropout to study the sensitivity of model to missing modality. In this setting, we empirically find that those methods with missing modality imputation usually shows higher tenacity in high proportion of missingness such as FlexMoE and our propose method. Although ShaSpec has also impute the missing modality by shared information, but it is weak on the backbone model and struggling to optimize with multi-task objectives. Our proposed method, on the other hand, shows superior performance across all setting in CMU-MOSEI and CMU-MOSI for all metrics.
Experiment Setting III: In Appendix E Table 11, Table 12, Table 13 and Table 14, we study which modality combination is significant and how is the tenacity of our proposed method to different distribution shift on missing modality. We find that single modality shows very similar performance during inference phrase in our proposed method and Flex-MoE across AUC and F1, indicating the effectiveness of missing modality imputation. In particular, our proposed also shows the highest resistance to missing modality across different combination and remains competitive performance no matter what modality is missing. In addition, the modality combination Audio-Video is the worse modality combination for all implemented models. The potential reason may be that the data interaction between those two modalities are weak or even opposite. We show some visualization evidence in Figure 4 that the embedding of those two modalities are mostly symmetry.
Table 3: Ablation I: Module dropout. CMU-MOSI adopt 50% missing in Experiment III
MIMIC-III CMU-MOSI
F1 AUC F1 AUC
w/o impute 43.79 $\pm$ 2.81 85.33 $\pm$ 1.22 42.51 $\pm$ 2.25 69.67 $\pm$ 1.27
w/o Conf 48.59 $\pm$ 1.97 85.91 $\pm$ 0.84 42.47 $\pm$ 0.94 68.77 $\pm$ 1.66
w/o impute & Conf 46.32 $\pm$ 1.66 85.15 $\pm$ 0.57 42.50 $\pm$ 0.63 66.73 $\pm$ 1.41
Conf-SMoE 53.44 $\pm$ 0.27 87.05 $\pm$ 0.58 43.89 $\pm$ 0.51 71.51 $\pm$ 0.53
4.5 Are Missing Modality Imputation and ConfNet Actually Useful?
We further implement Ablation Study I in Table 3 by dropping off the module we designed in proposed method and study which part of of model is actually useful. Specifically, we dropout the modality imputation module and keep the missing modality as zeros, we can see that F1 and AUC dropped off by 6.21% and 1.81% on MIMIC-III, 5.62% and 0.81% on CMU-MOSI. This performance drops is even worse when imputation and ConfNet gating mechanism is removed from proposed method, providing strong empirical evidence for effectiveness of our design. It is interesting that the w/o impute & Conf variant outperform w/o impute, indicating the model may only focus on existing modality instead of modality interaction.
@@ 4.6 Why Does Confidence-Guided Gating Improve Performance?
(a) Expert selection of softmax w/ $\mathcal{L}_{laod}$
@@
Table 4: Ablation II: Different router. CMU-MOSI adopt 50% missing setting in Experiment III
@@ 4.7 Complexity Analysis
Table 5: Complexity analysis on CMU-MOSI with 50% missing setting in Experiment III
Metrics ShaSpec TF LlMoE mmFormer FuseMoE FlexMoE ConfSMoE
MFLOPs 0.51 8.42 1016.15 1577.63 199.80 20.16 47.10
#Params 102,022 4,211,331 3,457,819 11,373,315 243,543,180 1,126,147 3,116,310
F1 score 36.39 $\pm$ 0.97 38.46 $\pm$ 1.09 41.54 $\pm$ 0.77 41.14 $\pm$ 0.31 40.90 $\pm$ 0.49 41.95 $\pm$ 0.30 43.89 $\pm$ 0.51
In Table 5, we compared the complexity between our proposed method and baseline models in terms of FLOPs and number of parameters. ShaSpec is the simplest models, but it turns out not comparable performance in many tasks. On the other hand, those MoE-based approach like FuseMoE and LlMoE are extremely complicated either in FLOPs and number of parameters while the performance is suboptimal compare to FlexMoE and our proposed method. mmFormer show very comparable results with our proposed method, however, the FLOPs and parameters are extremely larger than ConfSMoE, leading to poor training and inference efficiency. On the other hand, FlexMoE is 2 times less complicated than ConfSMoE while the performance is worse than LlMoE, mmFormer and ConfSMoE. By balancing complexity and performance from Table 5, it is clear that ConfSMoE is the best model to balance complexity and performance.
@@ 5 Related Works
While in many works considers that the interaction between multimodal data can bring comprehensive understanding for downstream task. Initial multimodal fusion approach usually incoperates neural kernel fusion[4, 24], early/late network fusion [32, 3, 11]. Approaches such as Tensor Fusion Network[37], Multimodal Transformer[28] and Multimodal Adaption Gate[25], LlMoE[22] all highlights the multimodal interactions are beneficial to downstream task. In medical domain, MedFuse [14] incorporated CXR and EHR modality to jointly fuse modality with early and later fusion in a single fusion network. MISTS [38] incorporated multi-time attention mechanism to encoder irregular time serires and adapts multimodal cross attention to fuse two different modality. [33] fused multimodal data in a attention gate and compute a replacement vector for modality fusion. However, these approaches are limited to full modality only and difficult to extend to increasing number of modality since the complexity can be significantly increased by the number of modality. Missing modality is realistic and practical problem in real life application such as medical diagnose. SMIL [20] proposed a Bayesian meta-learning solution with reconstruction network to impute the missing modality from in hidden state during fusion. [8, 26] propose generative adversarial networks to impute the missing modality by semi-supervised learning. ShaSpec [29] and DrFuse [34] designed multiple loss function to learn share and specific representation of different modality. However, the reconstruction network from SMIL and ShaSpec requires multiple additional auxiliary loss function, which occupies the great amount of gradient in optimization process instead of task specific loss. [27] In addition, recent studies like FuseMoE [12] respect the missingness of modalities and wish to assign each combination of multimodal data to different experts. Flex-MoE[35] further expand the advantages of FuseMoE and introduce a modality bank to impute the missing modality.

### v2

@@ 3.2 Missing Modality Modeling: Two-Stage Imputation
Inspired by the modality gap phenomenon observed in contrastive learning Liang et al. (2022), we hypothesize that a similar gap exists in our multimodal setting, and that it may even contribute positively to model performance, as illustrated in Appendix Figure 4. The presence of this modality gap suggests that imputing a missing modality using only shared information from the available modalities is inherently challenging, due to the distinct representational spaces each modality occupies. Ideally, the imputed modality should preserve the characteristics of its original distribution without being significantly biased by other modalities. To address this, we propose a Two-stage Imputation strategy to reconstruct the representative feature from the missing modality distribution itself, and further refine the imputed modality by adding up instance-specific feature from the available modalities.
Pre-Imputation: To retrieve modality-specific feature, given a sample $x_{i}$ with missing modality $M_{m}\in\mathbb{R}^{s\times d}$ and arbitrary available modality $M_{a}\in\mathbb{R}^{s\times d}$, where $s$ is the sequence length. Pre-imputation aims to obtain the modality-specific feature from other instances of $M_{m}$ distribution. One typical way is to take the mean feature distribution of $M_{m}$, but simply taking the mean may be overfitting to training set and all representation across different modality samples will be deterministic and uninformative. In contrast to mean representation, we wish to introduce some stochasticity during the pre-imputation as this stochasticity may prevent the representation to collapse to the same representation. Specifically, we generate an pre-imputed modality embedding $\bar{M}_{m}$ by sampling $n$ number of instances out of the modality pool $M_{m,i}=[M_{m,1},M_{m,2},\cdots,M_{m,n}]$.
$\bar{M}_{m}=\frac{1}{n}\sum^{n}_{i=1}\mathbf{1}M_{m,i}$ (6)
@@
Where $\mathbf{1}\in\{0,1\}^{n}$ is the selection vector for $M_{m}$ and $\mathbf{1}_{i}=1$ represents the $i$-th instance from modality pool is selected. $\bar{M}_{m}$ will be sent to ConfSMoE for learning the interaction across different modalities. We denoted $M_{m}^{*}$ as the pre-imputed modality representation after ConfSMoE and $M^{k}_{a}$ as the available modality representation after k-th expert.
@@
Post-Imputation: To refine $M^{*}_{m}$ by available modality, we proposed a post-modality imputation to further capture the interactions between available and missing modality. Unlike conventional approaches that solely rely on the shared representation from one network, our method leverages information from Top-K experts networks by selectively incorporating tokens. This allows the model to absorb nuanced, instance-specific features and interaction between modalities from different views of expert to available modality. We employ a sparse attention mechanism, under the assumption that only a subset of modality-specific information is meaningfully shared across $M^{k}_{a}$. By enforcing sparsity, the model is encouraged to attend to the most relevant parts of the available modalities, thereby reducing the risk of injecting modality-irrelevant or noisy information into the imputed representation. Recall from Eq. 1, Top-K MoE generate $K$ specialized representation from available modality $M_{a}^{k}\in\mathbb{R}^{s\times d}$, $k=0,1,\dots K$. Then, we concatenate all $M_{a}^{k}$ returned from $K$ expert to form opinion representation from different expert, denoted as $M_{a}^{*}\in\mathbb{R}^{(s\times K)\times d}$. We apply a sparse cross attention over $M_{m}^{*}$ and $M_{a}^{*}$ to select most relevant tokens to refine $M_{m}^{*}\in\mathbb{R}^{s\times d}$. The SparseCrossAttention takes $M_{m}^{*}$ as Query modality $\mathbf{Q}=\mathbf{W}_{q}M_{m}^{*}$, $M_{a}^{*}$ as Key modality $\mathbf{K}=\mathbf{W}_{k}M_{a}^{*}$ and Value modality $\mathbf{V}=\mathbf{W}_{v}M_{a}^{*}$, where $\mathbf{Q}\in\mathbb{R}^{s\times d}$ and $\mathbf{K},\mathbf{V}\in\mathbb{R}^{(s\times K)\times d}$. Note that $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$ are shared between all modalities. Then, we obtained the cross attention score $\mathbf{A}\in\mathbb{R}^{s\times(s\times K)}$ from the dot product of $\mathbf{Q}$ and $\mathbf{K}$. Then, we apply a binary mask $\mathbf{M}$ over the attention scores $\mathbf{A}$ to retain only the Top-T entry based on the rank of attention score (1 indicate presence) alongside the dimension of $\mathbf{K}$, yielding a sparse attention map $\mathbf{A}^{*}$. $T=\frac{s(|M|-1)}{B}$, where $|M|$ represents the number of modality in dataset and $B$ is a sparsity hyperparameter, we set $B$ as 4 by default.
$\mathbf{A}^{*}=\mathbf{M}\odot softmax(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}),\in\mathbb{R}^{s\times(s\times K)}$ (7)
Thus, we aggregate the value matrix $V$ using the sparse attention weights $A^{*}$ to obtain the output of $SparseCrossAttention$ (SCA). When multiple available modality $M_{a}^{*}$ are presented in the dataset, we compute their contributions individually and sum them to refine $M_{m}^{*}$:
$M^{*}_{m}=LayerNorm\left(M_{m}^{*}+\sum^{a}SCA(M_{m}^{*},M_{a}^{*},M_{a}^{*})\right)=LayerNorm\left(M_{m}^{*}+\sum^{a}\mathbf{A}^{*}\mathbf{V}\right)$ (8)
In addition, we expect this refinement to redistribute the representation of $M_{m}^{*}$ instead of generating large scale representation and lead to diverge training. A LayerNorm layer is applied to scale the feature map back to consistent distribution as other layers. With SCA, the Post-imputation serves as a refinement module to add up instance-specific feature on $M_{m}^{*}$, which each modality can contribute differently on different sub-segment representation of $M_{m}^{*}$. We visualize the attention weights after imputation in the Appendix. Figures 13, 14, and 15 illustrate these weights, showing that the query modality can dynamically adjust the significance of information by leveraging both expert specialization and shared cross-modal cues. The attention maps reveal distinct weight distributions across different modalities and experts.
@@ 4.1 Experiment Setting
To ensure a fair comparison with baseline methods, we adopt the best-performing hyperparameter settings reported in their respective papers and official implementations. Additionally, we standardize the hidden dimension to 128 across all models. To eliminate confounding effects introduced by varying encoder architectures, we use the same modality-specific encoders for all models, as differences in encoder design can significantly influence final performance. For the MIMIC-III and MIMIC-IV datasets, we employ a pretrained ClinicalBERT Alsentzer et al. (2019) for textual data, a one-layer patch embedding module for both irregular time series and ECG signals, and a pretrained DenseNet Cohen et al. (2022) from TorchXRayVision for chest X-ray (CXR) images. Experiment Setting I: Natural Missing Modality. This setting reflects real-world missingness patterns observed in clinical datasets. Experiment Setting II: Random Modality Dropout. A fixed percentage of modalities are randomly dropped for each instance, while ensuring that each instance retains at least one modality in both train and test set. Experiment Setting III: Asymmetric Modality Dropout. Half of the modalities are randomly dropped during training, while testing is performed with only one or two modalities present per instance, simulating domain shifts or deployment constraints. All experiments are 3-Fold cross-validate in different seeds for fair comparison.
@@ 4.2 Primary Results
Task Metric SMIL ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfSMoE-T ConfSMoE-E
@@
Experiment Setting II: In Appendix Table 9 and Table 10, we implement Random Modality Dropout to study the sensitivity of model to missing modality. In this setting, we empirically find that those methods with missing modality imputation usually shows higher tenacity in high proportion of missingness such as FlexMoE and our propose method. Although ShaSpec has also impute the missing modality by shared information, but it is weak on the backbone model and struggling to optimize with multi-task objectives. Our proposed method, on the other hand, shows superior performance across all setting in CMU-MOSEI and CMU-MOSI for all metrics.
Experiment Setting III: In Appendix Table 11, Table 12, Table 13 and Table 14, we study which modality combination is significant and how is the tenacity of our proposed method to different distribution shift on missing modality. We find that single modality shows very similar performance during inference phrase in our proposed method and Flex-MoE across AUC and F1, indicating the effectiveness of missing modality imputation. In particular, our proposed also shows the highest resistance to missing modality across different combination and remains competitive performance no matter what modality is missing. In addition, the modality combination Audio-Video is the worse modality combination for all implemented models. The potential reason may be that the data interaction between those two modalities are weak or even opposite. We show some visualization evidence in Figure 4 that the embedding of those two modalities are mostly symmetry.
@@ 4.3 Why Does Confidence-Guided Gating Improve Performance?
(a) Expert selection of softmax w/ $\mathcal{L}_{load}$
@@
Table 2: Ablation I: Different router. CMU-MOSI adopt 50% missing setting in Experiment III
@@
4.4 How ConfNet and Modality Imputation integrate together?
Table 3: Ablation II: Module dropout. CMU-MOSI adopt 50% missing in Experiment II
Variants MIMIC-III CMU-MOSI
F1 AUC F1 AUC
w/o Conf 48.59 $\pm$ 1.97 85.91 $\pm$ 0.84 42.47 $\pm$ 0.94 68.77 $\pm$ 1.66
w/o impute 49.59 $\pm$ 0.29 86.31 $\pm$ 0.12 42.46 $\pm$ 2.29 68.44 $\pm$ 1.82
w/o post-impute 49.86 $\pm$ 1.89 86.53 $\pm$ 0.10 43.14 $\pm$ 1.32 69.13 $\pm$ 1.45
w/o impute & Conf 46.32 $\pm$ 1.66 85.15 $\pm$ 0.57 41.98 $\pm$ 0.74 68.48 $\pm$ 1.90
ConfSMoE-T 53.44 $\pm$ 0.27 87.05 $\pm$ 0.58 44.34 $\pm$ 0.77 70.41 $\pm$ 1.13
We implement the experiment in Table 3 by removing ConfNet and the two-stage imputation module gradually. This experiment demonstrates that both novel modules, ConfNet and the two-stage imputation, are effective on their own, while their integration provides strong robustness against the missing modality challenge. Specifically, ConfNet captures informative features from the imputed modality, and the adaptive two-stage imputation operation reconstructs high-quality features, thereby complementing each other to achieve robust performance. Specifically, we dropped the entire imputation module and keep the missing modality as zeros (w/o impute), we can see that F1 and AUC dropped off by 3.85 and 0.74 on MIMIC-III, 1.88 and 1.97 on CMU-MOSI. This performance drops is even worse when imputation and ConfNet gating mechanism is removed from proposed method, providing strong empirical evidence for effectiveness of our design. We further remove the post-imputation (w/o post-impute) and observe that ConfNet can still take advantage of modality-specific feature for their downstream task, showing robustness against missing modality.
@@ 4.5 Complexity Analysis
In Table 4, we control the number of parameters for both baseline and proposed methods to compare model complexity and performance. Although ShaSpec and TF achieve lower FLOPs for the same parameter amount, their performance remains less competitive than most Transformer- and MoE-based models, except FuseMoE. Among MoE-based models, ConfSMoE-T delivers the best performance while maintaining similar MFLOPs and parameter counts to the second-best baseline. Under comparable parameter settings, LIMoE and mmFormer fail to keep FLOPs as low as ConfSMoE-T and yield suboptimal results. Moreover, LIMoE is particularly inefficient, as its MFLOPs are substantially higher than those of other models with similar parameter sizes. Overall, ConfSMoE-T achieves the best balance between efficiency and performance among MoE-based baselines.
Table 4: Complexity analysis I: Small Scale
Metrics ShaSpec TF LlMoE mmFormer FuseMoE FlexMoE ConfSMoE-T
MFLOPs 21,73 4.20 1245.03 237.39 143.93 40.32 39,62
#Params 2,177,406 2,101,571 2,130,227 2,378,371 20,971,020 2,184,963 2,184,730
F1 score 39.12 $\pm$ 1.21 35.81 $\pm$ 0.85 41.11 $\pm$ 0.90 41.14 $\pm$ 0.31 36.52 $\pm$ 0.84 42.12 $\pm$ 1.18 44.63 $\pm$ 1.65
@@ 5 Related Works
While in many works considers that the interaction between multimodal data can bring comprehensive understanding for downstream task. Initial multimodal fusion approach usually incoperates neural kernel fusionBucak et al. (2013); Poria et al. (2015), early/late network fusion Xu et al. (2023); Baltrušaitis et al. (2018); Guo et al. (2019). Approaches such as Tensor Fusion NetworkZadeh et al. (2017), Multimodal TransformerTsai et al. (2019) and Multimodal Adaption GateRahman et al. (2020), LlMoEMustafa et al. (2022) all highlights the multimodal interactions are beneficial to downstream task. In medical domain, MedFuse Hayat et al. (2022) incorporated CXR and EHR modality to jointly fuse modality with early and later fusion in a single fusion network. MISTS Zhang et al. (2023) incorporated multi-time attention mechanism to encoder irregular time serires and adapts multimodal cross attention to fuse two different modality. Yang and Wu (2021) fused multimodal data in a attention gate and compute a replacement vector for modality fusion. However, these approaches are limited to full modality only and difficult to extend to increasing number of modality since the complexity can be significantly increased by the number of modality. Missing modality is realistic and practical problem in real life application such as medical diagnose. SMIL Ma et al. (2021) proposed a Bayesian meta-learning solution with reconstruction network to impute the missing modality from in hidden state during fusion. Du et al. (2018); Shang et al. (2017) propose generative adversarial networks to impute the missing modality by semi-supervised learning. ShaSpec Wang et al. (2023) and DrFuse Yao et al. (2024) designed multiple loss function to learn share and specific representation of different modality. However, the reconstruction network from SMIL and ShaSpec requires multiple additional auxiliary loss function, which occupies the great amount of gradient in optimization process instead of task specific loss. Sun et al. (2024) In addition, recent studies like FuseMoE Han et al. (2024) respect the missingness of modalities and wish to assign each combination of multimodal data to different experts. Flex-MoEYun et al. (2024) further expand the advantages of FuseMoE and introduce a modality bank to impute the missing modality.